# ECG-Byte: A Tokenizer for End-to-End Generative Electrocardiogram Language Modeling

### v1

@@ Abstract
Large Language Models (LLMs) have shown remarkable adaptability across domains beyond text, specifically electrocardiograms (ECGs). More specifically, there is a growing body of work exploring the task of generating text from a multi-channeled ECG and corresponding textual prompt. Current approaches typically involve pretraining an ECG-specific encoder with a self-supervised learning (SSL) objective and using the features output by the pretrained encoder to finetune a LLM for natural language generation (NLG). However, these methods are limited by 1) inefficiency from two-stage training and 2) interpretability challenges with encoder-generated features. To address these limitations, we introduce ECG-Byte, an adapted byte pair encoding (BPE) tokenizer pipeline for autoregressive language modeling of ECGs. This approach compresses and encodes ECG signals into tokens, enabling end-to-end LLM training by combining ECG and text tokens directly, while being much more interpretable since the ECG tokens can be directly mapped back to the original signal. Using ECG-Byte, we achieve competitive performance in NLG tasks in only half the time and $\sim$48of the data required by two-stage approaches.
Data and Code Availability
This paper uses the ECG-Chat pretraining (Zhao et al., 2024) and ECG-QA datasets (Oh et al., 2023), which were both created from the MIMIC-IV ECG (Johnson et al., 2023) and PTB-XL datasets (Wagner et al., 2020). More details about the datasets are provided in Section 3. The ECG-Chat pretraining, ECG-QA, MIMIC-IV ECG, and PTB-XL datasets are all freely available on https://github.com/YubaoZhao/ECG-Chat, https://github.com/Jwoo5/ecg-qa, and https://physionet.org/ respectively. Lastly, we have released the code at the following link: https://github.com/willxxy/ECG-Byte.
Institutional Review Board (IRB)
All datasets used in this study are directly taken from the publicly available, de-identified MIMIC-IV ECG (Johnson et al., 2023) and PTB-XL (Wagner et al., 2020) datasets, thus not requiring IRB approval.
@@ 1 Introduction
Figure 1: Comparison of traditional methods and our method on ECG language modeling. Traditional methods comprises 2 training stages. The first stage aims to learn a good representation of the 12-lead ECG by training an ECG-specific encoder with a self-supervised learning objective via a combinatorial or individual usage of contrastive learning ($L_{CL}$) on the hidden states of the ECG $z_{e}$ and text $z_{t}$ or masked image modeling ($L_{MIM}$). The second stage applies the trained encoder to an ECG, which is input alongside a text prompt to a LLM for generation. In contrast, our method is end-to-end and can directly train a LLM for generation utilizing ECG-Byte.
@@
The aforementioned facts calls attention to the need for accessible, accurate, and efficient automation of ECG analysis through deep learning. Deep learning has reached expert level performance in certain tasks for CVD detection using ECGs (Rajpurkar et al., 2017; Hannun et al., 2019; Qiu et al., 2023b). However, most previous works in this domain have succumbed to a crude classification of hard CVD labels (Choi et al., 2023; Nonaka and Seita, 2020; Martin et al., 2021; Strodthoff et al., 2021). A problem with this approach is that ECGs often do not exclusively fall into one diagnostic category, instead, there may be many soft labels annotated by expert physicians and the accumulation of these soft labels allow a more detailed, nuanced, and clinically useful interpretation of the ECG.
@@
The recent onset of Large Language Models (LLMs) provides an opportunity to take a softer, generative, and consequently more flexible approach to ECG analysis. There are some recent works that have taken advantage of this approach to ECG analysis (Qiu et al., 2023a; Tang et al., 2024b; Wan et al., 2024; Fu et al., 2024; Zhao et al., 2024). A commonality among these works is that they treat multi-channel ECGs similarly to images; they first pretrain an encoder specifically for ECGs with some self-supervised learning (SSL) objective, then apply the learned features to finetune a LLM for natural language generation (NLG). However, we observed 2 limitations with this approach: 1) inefficiencies from two-stage training and 2) interpretability challenges with encoder generated features. Pretraining a good ECG specific encoder can be a significant computational burden due to large datasets, model size, and long training times. Additionally, the latent feature vector output by the ECG encoder cannot be mapped back to the signal, making interpretability difficult when utilizing this feature vector for downstream tasks.
@@
In this study, we introduce ECG-Byte, an adapted byte pair encoding (BPE) (Gage, 1994) tokenizer designed for end-to-end training for generative ECG language modeling. Inspired by prior works demonstrating the effectiveness of creating discrete tokens from continuous values (Chen et al., 2022; Han et al., 2024), we leverage quantization to represent amplitude ranges as discrete symbols. Using discrete symbols, we obtain string representations of the ECGs to train ECG-Byte. We then apply ECG-Byte to directly finetune a LLM for conditional autoregressive NLG, where the text output is conditioned on the text prompt and tokenized signal. We found that our end-to-end training approach is competitive in conditional NLG with only half the time and $\sim$48of the data required by 2-stage pretraining approaches. Additionally, since the encoded ECG can be reverse transformed back to the original signal, we can interpret token-level attention-based visualizations, whereas it would be impossible to reverse the hidden latent feature vector outputs by a pretrained encoder.
Our contributions are summarized as the following:
  1. 1.
    We present ECG-Byte, an adapted BPE tokenizer for end-to-end training on autoregressive, conditional NLG.
  2. 2.
@@
    We empirically show the efficiency of our method and present competitive performance compared to conventional 2-stage pretraining approaches, proposing a new paradigm for conditional NLG with ECGs.
  3. 3.
    We conduct an interpretability study on both ECG-Byte and the LLM by respectively observing how ECG-Byte is merging ECG signals and by utilizing attention visualizations to observe how the LLM is processing ECGs and text.
@@ 2.1 Deep learning for ECGs
There has been a plethora of works utilizing deep learning for processing ECGs for classification (Rajpurkar et al., 2017; Hannun et al., 2019; Choi et al., 2023; Nonaka and Seita, 2020; Martin et al., 2021; Strodthoff et al., 2021). Most of these works utilize either convolutional neural networks (CNNs) (Rajpurkar et al., 2017) or transformers (Choi et al., 2023) and exhibit excellent performance at classification tasks. There have also been some efforts to frame classification as a retrieval task in order to recover cases similar to the given ECG.(Tang et al., 2024b; Qiu et al., 2023b). However, the retrieval approach may struggle with rare or unique ECG patterns that lack good matches in the existing database, potentially leading to missed or inaccurate diagnoses for unusual cases. Additionally, both classification and retrieval tasks may be crude formulations of processing ECGs, since ECGs typically exhibit many characteristics of overlapping CVDs.
@@ 2.2 Large Language Models for ECGs
Generative Large Language Models (LLMs) have given the opportunity to take a softer and more clinically similar approach in processing ECGs by generating physician-vetted clinical statements (Qiu et al., 2023a; Tang et al., 2024b; Wan et al., 2024; Fu et al., 2024; Zhao et al., 2024). The representation of ECG data has been largely considered in previous works which feed the raw signal into an encoder to obtain a latent representation of the ECG data, which then serves as input into a LLM. (Zhao et al., 2024; Wan et al., 2024; Tang et al., 2024b; Choi et al., 2023). In order to get good latent representations of the ECGs, an encoder is first pretrained on a self-supervised learning (SSL) objective (e.g., contrastive learning, masked language/image modeling). Although model performance in terms of label classification has been excellent when using these approaches, we want to be able to generate soft labels akin to clinical notation since ECGs often have overlapping and non-mutually exclusive descriptors. There have been some efforts in this direction (Qiu et al., 2023a; Tang et al., 2024b; Wan et al., 2024; Fu et al., 2024; Zhao et al., 2024), however, they suffer in efficiency (i.e., requires 2 stages of training). In our work, we challenge this 2-stage pretraining approach by transforming the ECG into tokens using ECG-Byte and directly training a LLM for NLG.
@@ 2.3 Byte Pair Encoding for Domains Outside of Language
The Byte Pair Encoding (BPE) algorithm was first introduced by Gage (1994) for data compression. It was later adapted to the natural language processing (NLP) domain (Sennrich et al., 2016) and has been the favored approach to tokenization for the most popular language models (Grattafiori et al., 2024; Brown et al., 2020) due to its efficiency and robustness to rare words. Byte pair encoding has seen success in representing modalities outside of language, including molecular graphs (Shen and Póczos, 2024), electroencephalogram (EEG) (Klymenko et al., 2023), and more generally, physiological signals.(Tavabi and Lerman, 2021). However, in theses cases the byte pair encoded representations are simply used for classification. Most recently, Tahery et al. (2024) leveraged quantization and BPE to compress ECG signals and pass as inputs to a BERT (Devlin et al., 2019) model for SSL. However, in their work, they only use this representation for classification. As previously mentioned, we believe classification alone may limit aspects of ECG interpretation, thus we utilize these representations for generative diagnosis. Additionally, previous works utilized a pre-existing BPE tokenizer based on SentencePiece (Kudo and Richardson, 2018), and do not conduct further analysis of how the BPE algorithm is merging the ECGs.
@@ 3.1 Dataset and Preprocessing
In this study, we use variants of the MIMIC-IV ECG (Gow et al., 2023) and PTB-XL datasets (Wagner et al., 2020) for NLG. We use MIMIC-IV ECG pretraining curated by Zhao et al. (2024) that contains question prompts generated by GPT-4o alongside the ECG and clinical notes. Additionally, we use the ECG-QA dataset (Oh et al., 2023), a dataset that uses the ChatGPT API to generate naturalistic, clinically relevant question and answer pairs about the ECG signals from the MIMIC-IV ECG and PTB-XL datasets. The baselines we compare our results with all utilize the single-verify, single-choose, and single-query categorized questions from the ECG-QA dataset. single-verify corresponds to yes or no questions, single-choose corresponds to where a selected answer is made from two given options, and single-query consists of open-ended questions. The ECG signals collected from both both datasets (i.e., MIMIC-IV ECG and PTB-XL) are sampled at 500 Hz for 10 seconds, resulting in a 5000 length, 12 lead ECG.
@@
We preprocess all datasets used in the study in the same manner to maintain consistency. We first convert the ordering of the leads for the MIMIC-IV ECG dataset (i.e., [‘I’, ‘II’, ‘III’, ‘aVR’, ‘aVF’, ‘aVL’, ‘V1’, ‘V2’, ‘V3’, ‘V4’, ‘V5’, ‘V6’])to the PTB-XL dataset format (i.e., [‘I’, ‘II’, ‘III’, ‘aVL’, ‘aVR’, ‘aVF’, ‘V1’, ‘V2’, ‘V3’, ‘V4’, ‘V5’, ‘V6’]). We then use a notch filter at 50 Hz and 60 Hz to remove powerline interference. Each frequency is targeted with a quality factor of 30 to minimize distortion, and filtering is applied bidirectionally to prevent phase shifts. Next, a fourth-order Butterworth bandpass filter with a range of 0.5–100 Hz isolates relevant ECG components while attenuating high-frequency noise and low-frequency drift. To address baseline wander caused by respiratory or movement artifacts, we bidirectionally apply a fourth-order Butterworth highpass filter with a cutoff of 0.05 Hz. After filtering, we apply wavelet denoising to further reduce noise. Using the Daubechies-6 (db6) wavelet at level 4, we decompose each ECG signal into wavelet coefficients. A soft threshold, based on the median absolute deviation of the detail coefficients, is applied to each coefficient level to suppress noise, ensuring values near zero are excluded from reconstruction. Since 250 Hz is a generally accepted sampling frequency adequate for heartbeat analysis (Kwon et al., 2018), we downsample the 500 Hz sampling frequency to 250. We then segment the 10 second signal to non-overlapping windows of 2 seconds, and use each 12 lead 2 second segment of the ECG signal as input to the model. However, for training the tokenizer, we did not want to introduce this discontinuity across the full 10 seconds. Thus, we utilize the unsegmented, 10 second ECG signal for training ECG-Byte. Lastly, during the unsegmented preprocessing pipeline, we record the global 1st and 99th percentiles out of 300,000 samples to utilize in our later steps of training ECG-Byte for normalization.
@@ 3.2 ECG as Bytes
Following established practices in NLP (Dagan et al., 2024), we train ECG-Byte on a representative subset of the total dataset, selected using stratified sampling based on morphological clustering. To extract features from each unsegmented ECG, we compute statistical measures, frequency and time domain features, morphological characteristics, and wavelet coefficients. Principal Component Analysis (PCA) (Wold et al., 1987) is applied for dimensionality reduction, retaining 95% of the variance, followed by feature scaling. The optimal number of clusters is determined using the Elbow Method and Silhouette Analysis (Rousseeuw, 1987), with the smaller result chosen. K-means clustering (MacQueen, 1967) is then applied to the scaled PCA-transformed features. If K-means fails to yield distinct clusters, DBSCAN (Ester et al., 1996) is used as a fallback. Stratified sampling is performed by randomly selecting ECGs from each cluster in proportion to its size, resulting in a total sample of 200,000 ECGs.
@@
where $\epsilon_{1}=0.5$ is a constant to make up for the sampled percentiles and $\epsilon_{2}=10^{-6}$ is a small constant added to prevent division by zero. This transformation shifts and scales $X$ so that the normalized values fall within the range [0, 1]. We then apply clipping to ensure that values remain strictly within this range:
$X_{\text{clipped}}=\text{clip}(X_{\text{norm}},0,1)$ (2)
Inspired by previous works (Klymenko et al., 2023; Tavabi and Lerman, 2021; Chen et al., 2022; Han et al., 2024), we quantize $X_{\text{clipped}}$ into discrete levels for symbolic representation. Let $\mathcal{A}$ be the set of 26 symbols, corresponding to the lowercased letters in the English alphabet, $\mathcal{A}=\{\text{a},\text{b},\ldots,\text{z}\}$. The alphabet size $|\mathcal{A}|=26$ defines the number of discrete levels. We scale and floor $X_{\text{clipped}}$ to integer values, then take the minimum between the floored value and the maximum number of bins as the following:
$X_{\text{quant}}=\text{min}(\lfloor X_{\text{clipped}}\times|\mathcal{A}|\rfloor,(|\mathcal{A}|-1))$ (3)
Finally, each integer value in $X_{\text{quant}}$ is mapped to a corresponding symbol in $\mathcal{A}$ to yield the symbolic signal, which serves as a discrete representation of the ECG. After transforming each ECG signal instance into its symbolic form, we first flatten each symbolic ECG instance $X_{\text{quant}}^{(i)}$ into a 1-dimensional sequence of symbols:
$X_{\text{symb}}^{(i)}=\text{flatten}(X_{\text{quant}}^{(i)}),\quad X_{\text{symb}}^{(i)}\in\mathcal{A}^{CT}$ (4)
where $i$ indexes over all instances in the dataset, and $X_{\text{symb}}^{(i)}$ is the flattened sequence of symbols of length $C\cdot T$. Next, we concatenate all flattened instances $X_{\text{symb}}^{(1)},X_{\text{symb}}^{(2)},\ldots,X_{\text{symb}}^{(N)}$ across the entire dataset to form a single, long symbolic sequence:
$\mathbf{X}_{\text{concat}}=X_{\text{symb}}^{(1)}\|X_{\text{symb}}^{(2)}\|\cdots\|X_{\text{symb}}^{(N)},\quad\mathbf{X}_{\text{concat}}\in\mathcal{A}^{NCT}$ (5)
[…]
@@
After obtaining the string representation $\mathbf{X}_{\text{concat}}$ of the ECG dataset, we train ECG-Byte to compress the discretized ECG signals by iteratively merging the most frequent byte pairs into single tokens, following the BPE algorithm. The process starts by converting $\mathbf{X}_{\text{concat}}$ into an ID vector of 8-bit unsigned integers and initializing a vocabulary map (vocab) for string representations of bytes and a vocab_tokens map to encode bytes as singleton lists. IDs and vocab are initialized to cover the full byte range (0–255), mapping symbols in $\mathcal{A}$ to ASCII values (97–122), while reserving other byte values for unknown bytes. As merging proceeds, new tokens are assigned unique integer IDs starting from 256, acting as abstract labels for progressively larger token units. For each merge iteration, ECG-Byte calculates adjacent byte pair frequencies using a parallelized get_stats function, efficiently aggregating counts via a fold-and-reduce strategy. The most frequent pair is identified as the ”best pair” to merge, and the merge function replaces occurrences of this pair in the ID vector with a new token ID, extending the vocabulary and updating vocab_tokens accordingly. This process repeats until the specified number of merges is reached or no pairs remain. The output includes the encoded ID vector, the extended vocabulary map, and a history of merge operations. Existing tokenizers, such as SentencePiece (Kudo and Richardson, 2018) or HuggingFace (Wolf et al., 2020), were not used due to their complexity and integration issues, which hindered interpretability. ECG-Byte, implemented in Rust for speed, provides a lightweight, flexible framework for representing ECG signals as discrete tokens while drawing inspiration from HuggingFace’s tokenizer (Wolf et al., 2020). Detailed pseudocode for the main training pipeline is provided in Algorithm 1, with merge and get_stats functions detailed in Appendix A.1.

### v2

@@ Abstract
Large Language Models (LLMs) have demonstrated exceptional versatility across domains, including applications to electrocardiograms (ECGs). A growing body of work focuses on generating text from multi-channeled ECG signals and corresponding textual prompts. Existing approaches often involve a two-stage process: pretraining an ECG-specific encoder with a self-supervised learning (SSL) objective, followed by finetuning an LLM for natural language generation (NLG) using encoder-derived features. However, these methods face two key limitations: inefficiency due to multi-stage training and challenges in interpreting encoder-generated features. To overcome these issues, we propose ECG-Byte, an adapted byte pair encoding (BPE) tokenizer pipeline for autoregressive language modeling of ECGs. ECG-Byte compresses and encodes ECG signals into tokens, enabling direct end-to-end LLM training by combining ECG and text tokens. This approach enhances interpretability, as ECG tokens can be directly mapped back to the original signals. Leveraging ECG-Byte, we achieve competitive NLG performance while training 3 times faster and using just 48% of the data required by traditional two-stage methods. All code is available at github.com/willxxy/ECG-Byte.
@@ 1 Introduction
Figure 1: Comparison of traditional and our approach for ECG language modeling. Traditional methods follow a two-stage process: (i) training a 12-lead ECG encoder with self-supervised objectives—contrastive learning ($L_{CL}$ between ECG ($z_{e}$) and the textual diagnostic report ($z_{o}$)) and/or masked image modeling ($L_{MIM}$)—to learn robust latent features; and (ii) mapping these ECG features ($z^{\prime}_{e}$) to a shared representation space via a learnable projection layer, then concatenating them with textual prompt vectors ($z_{t}$) as input to a large language model (LLM) for generation. In contrast, our method trains an LLM end-to-end using ECG-Byte as an ECG tokenizer.
@@
The aforementioned facts calls attention to the need for accessible, accurate, and efficient automation of ECG analysis through deep learning. Deep learning has reached expert level performance in certain tasks for CVD detection using ECGs (Rajpurkar et al., 2017; Hannun et al., 2019). However, most previous works have succumbed to a crude classification of hard CVD labels (Nonaka and Seita, 2020; Martin et al., 2021; Strodthoff et al., 2021). A problem with this approach is that ECGs often do not exclusively fall into one diagnostic category, instead, there may be many soft labels annotated by expert physicians and the accumulation of these soft labels allow a more detailed, nuanced, and clinically useful interpretation of the ECG (Singstad and Muten, 2022; Yoo et al., 2021).
@@
The recent emergence of Large Language Models (LLMs) enables a generative, flexible approach to ECG analysis. Recent works have adopted this method (Tang et al., 2024b; Zhao et al., 2024), treating multi-channel ECGs as images by first pretraining an ECG-specific encoder with a self-supervised learning (SSL) objective and then finetuning an LLM for natural language generation (NLG). However, this two-stage process has two key drawbacks: (1) training inefficiencies, as pretraining an effective encoder demands significant computational resources due to large datasets, model sizes, and long training times, and (2) interpretability challenges, since the encoder’s latent feature vectors cannot be mapped back to the original signal.
@@
In this study, we introduce ECG-Byte, a byte pair encoding (BPE) tokenizer (Gage, 1994) designed for end-to-end training of generative ECG language models. Building on prior work that converts continuous values to discrete tokens (Chen et al., 2022; Han et al., 2024), we use quantization to represent amplitude ranges as discrete symbols and obtain string representations of ECGs. We then directly finetune an LLM for autoregressive natural language generation (NLG), conditioning text output on both the text prompt and tokenized signal. Our end-to-end approach is competitive for conditional NLG, with 3 times faster total training time and approximately 48% of the data compared to two-stage pretraining methods. Moreover, since the tokenized ECG can be directly traced back to the original signal, token-level attention-based visualizations become interpretable.
Our contributions are summarized below:
1. We introduce ECG-Byte, a tokenizer that compresses discretized symbolic representations of ECGs using BPE, enabling direct end-to-end training without the need for an ECG-specific encoder.
@@
2. We empirically show the efficiency of our method and present competitive performances against conventional two-stage pretraining approaches, proposing a new paradigm for conditional NLG with ECGs.
3. We perform an interpretability study on ECG-Byte and the LLM, examining how ECG-Byte merges ECG signals and using attention visualizations to analyze how the LLM processes ECGs and text.
Generalizable Insights about Machine Learning in the Context of Healthcare
The motivation and generalizable insights of ECG-Byte derive from Vision-Language Models (VLMs) (Liu et al., 2023, 2024b), where image encoders are pretrained on large-scale data and later applied for visual understanding tasks. While recent ECG research has focused on developing specialized encoders for classification (Na et al., 2024; Liu et al., 2024a; Kiyasseh et al., 2021) and some works like ECG-Chat (Zhao et al., 2024) have shown potential for using these encoders in NLG, ECG-Byte demonstrates that a simpler approach can achieve competitive results. Our key insight is that compressing ECG data via a rule-based method (BPE tokenization) eliminates the need for complex encoder training pipelines that often involve additional overhead due to the size of the encoder, multiple loss objectives and carefully guided feature extraction (Gopal et al., 2021; Oh et al., 2022). Our analysis reveals that ECG-Byte effectively preserves clinically relevant signal components (P wave, QRS complex, T wave) within its tokenization scheme. Beyond cardiovascular applications, this successful adaptation of BPE for time series data suggests a promising approach for processing other physiological signals such as electroencephalograms (EEGs), photoplethysmograms (PPGs), or electromyograms (EMGs) in generative tasks with LLMs.
@@ 2.1 Deep learning for ECGs
There has been a plethora of works utilizing deep learning for processing ECGs for classification (Rajpurkar et al., 2017; Hannun et al., 2019; Nonaka and Seita, 2020; Martin et al., 2021; Strodthoff et al., 2021). Most of these works utilize either convolutional neural networks (CNNs) (Rajpurkar et al., 2017) or transformers (Choi et al., 2023) and exhibit excellent performance at classification tasks. There have also been some efforts to frame classification as a retrieval task in order to recover cases similar to the given ECG (Tang et al., 2024b; Qiu et al., 2023b). However, the retrieval approach may struggle with rare or unique ECG patterns that lack good matches in the existing database. Additionally, both classification and retrieval tasks may be crude formulations of processing ECGs, since ECGs typically exhibit many characteristics of overlapping CVDs.
@@ 2.2 Large Language Models for ECGs
Generative Large Language Models (LLMs) have given the opportunity to take a softer and more clinically similar approach in processing ECGs by generating physician-vetted clinical statements (Qiu et al., 2023a; Tang et al., 2024b; Wan et al., 2024; Fu et al., 2024; Zhao et al., 2024). Previous works have largely focused on representing ECG data by feeding the raw signal into a neural network encoder to output a latent representation, which is then used as input for an LLM (Zhao et al., 2024; Wan et al., 2024; Tang et al., 2024b). Recent studies have explored using printed ECGs as images for natural language generation (NLG) (Liu et al., 2024c; Khunte et al., 2024). However, our study focuses solely on methods that utilize the direct ECG signal. In order to obtain robust latent representations of the ECG signal, an encoder is first trained on a self-supervised learning (SSL) objective (e.g., contrastive learning, masked language/image modeling). Although model performance in terms of label classification has been excellent when using these approaches, we want to be able to generate soft labels akin to clinical notation since ECGs often have overlapping and non-mutually exclusive descriptors. While some efforts have been made in this direction (Tang et al., 2024b; Zhao et al., 2024), they face challenges in efficiency (requiring two stages of training) and interpretability (as the latent feature vectors from the ECG encoder are not interpretable). In our work, we challenge these two-stage pretraining approaches by transforming the ECG into tokens using ECG-Byte and directly training an LLM for NLG.
@@ 2.3 Byte Pair Encoding for Domains Outside of Language
The Byte Pair Encoding (BPE) algorithm was introduced by Gage (1994) for data compression and later adapted for natural language processing (NLP) (Sennrich et al., 2016). It is favored for tokenization in popular language models (Grattafiori et al., 2024; Brown et al., 2020) due to its efficiency and robustness to rare words. Beyond language, BPE has been applied to modalities such as molecular graphs (Shen and Póczos, 2024), electroencephalograms (EEG) (Klymenko et al., 2023), and other physiological signals (Tavabi and Lerman, 2021) for classification. Most recently, Tahery et al. (2024) used quantization and BPE to compress ECG signals as inputs to a BERT (Devlin et al., 2019) model for self-supervised learning, though only for classification. We argue that classification alone may limit ECG interpretation and therefore leverage these representations for generative diagnosis. Additionally, previous works employ a pre-existing BPE tokenizer (Tahery et al., 2024) based on SentencePiece (Kudo and Richardson, 2018) without further analyzing how the BPE algorithm merges ECGs. Inspired by HuggingFace (Wolf et al., 2020), we develop ECG-Byte, a custom Rust-based BPE tokenizer for ECGs. We conduct an extensive analysis of merged token usage, the distribution of tokenized ECG lengths, and provide visualizations mapping compressed tokens to the original ECG signal.
@@ 3.1 Dataset and Preprocessing
In this study, we use variants of the MIMIC-IV ECG (Gow et al., 2023) and PTB-XL datasets (Wagner et al., 2020) for NLG. We use MIMIC-IV ECG pretraining curated by Zhao et al. (2024) that contains question prompts generated by GPT-4o alongside the ECG and clinical notes. Additionally, we use the ECG-QA dataset (Oh et al., 2023), a dataset that uses the ChatGPT API to generate naturalistic, clinically relevant question and answer pairs about the ECG signals from the MIMIC-IV ECG and PTB-XL datasets. The baselines we compare our results with all utilize the single-verify, single-choose, and single-query categorized questions from the ECG-QA dataset. The ECG signals collected from both datasets (i.e., MIMIC-IV ECG and PTB-XL) are sampled at 500 Hz for 10 seconds, resulting in a 5000 length, 12 lead ECG.
@@
All datasets are preprocessed uniformly. For the MIMIC-IV ECG, we first reorder leads to match the PTB-XL format (from [I, II, III, aVR, aVF, aVL, V1–V6] to [I, II, III, aVL, aVR, aVF, V1–V6]). Powerline interference is removed using bidirectional notch filters at 50 Hz and 60 Hz (Q=30). A fourth-order Butterworth bandpass filter (0.5–100 Hz) isolates ECG components, while a bidirectional fourth-order highpass filter (cutoff 0.05 Hz) mitigates baseline wander. We then apply Daubechies-6 (db6) wavelet denoising at level 4, using a soft threshold based on the median absolute deviation of the detail coefficients. The signal is downsampled from 500 Hz to 250 Hz and segmented into non-overlapping 2-second windows for model input—except during tokenizer training, where the full 10-second signal is retained to avoid discontinuity. Finally, global 1st and 99th percentiles (from 300,000 samples) are recorded for later normalization in training ECG-Byte.
@@ 3.2 ECG as Bytes
Following established practices in NLP (Dagan et al., 2024), we train ECG-Byte on a representative subset of the total dataset, selected using stratified sampling based on morphological clustering. To extract features from each unsegmented ECG, we compute statistical measures, frequency and time domain features, morphological characteristics, and wavelet coefficients. Principal Component Analysis (PCA) (Wold et al., 1987) is applied for dimensionality reduction, retaining 95% of the variance, followed by feature scaling. The optimal number of clusters is determined using the Elbow Method and Silhouette Analysis (Rousseeuw, 1987), with the smaller result chosen. K-means clustering (MacQueen, 1967) is then applied to the scaled PCA-transformed features. If K-means fails to yield distinct clusters, DBSCAN (Ester et al., 1996) is used as a fallback. Stratified sampling is performed by randomly selecting ECGs from each cluster in proportion to its size, resulting in a total sample of 200,000 ECGs for training ECG-Byte.
@@
where $\epsilon_{1}=0.5$ is a constant to make up for the sampled percentiles and $\epsilon_{2}=10^{-6}$ is a small constant added to prevent division by zero. This transformation shifts and scales $X$ so that the normalized values fall within the range [0, 1]. We then apply clipping to ensure that values remain strictly within this range. Inspired by previous works (Klymenko et al., 2023; Tavabi and Lerman, 2021; Chen et al., 2022; Han et al., 2024), we quantize $X_{\text{norm}}$ into discrete levels for a symbolic representation. Let $\mathcal{A}$ be the set of 26 symbols, corresponding to the lowercased letters in the English alphabet, $\mathcal{A}=\{\text{a},\text{b},\ldots,\text{z}\}$. The alphabet size $|\mathcal{A}|=26$ defines the number of discrete levels. We scale and floor $X_{\text{norm}}$ to integer values, then take the minimum between the floored value and the maximum number of bins as the following:
$X_{\text{quant}}=\text{min}(\lfloor X_{\text{norm}}\times|\mathcal{A}|\rfloor,(|\mathcal{A}|-1))$ (2)
Finally, each integer value in $X_{\text{quant}}$ is mapped to a corresponding symbol in $\mathcal{A}$ to yield the symbolic signal, which serves as a discrete representation of the ECG. After transforming each ECG signal instance into its symbolic form, we first flatten each symbolic ECG instance $X_{\text{quant}}^{(i)}$ into a 1-dimensional sequence of symbols $X_{\text{symb}}^{(i)}$, where $X_{\text{symb}}^{(i)}\in\mathcal{A}^{CT}$, $i$ indexes over all instances in the dataset, and $X_{\text{symb}}^{(i)}$ is the flattened sequence of symbols of length $C\cdot T$. Next, we concatenate all flattened instances $X_{\text{symb}}^{(1)},X_{\text{symb}}^{(2)},\ldots,X_{\text{symb}}^{(N)}$ across the entire dataset to form a single, long symbolic sequence $X_{\text{concat}}$, where $X_{\text{concat}}\in\mathcal{A}^{NCT}$, and $N$ is the total number of instances in the dataset. The concatenated symbolic sequence $X_{\text{concat}}$ of length $N\cdot C\cdot T$ is then used to train ECG-Byte.
@@
After obtaining the string representation $\mathbf{X}_{\text{concat}}$ of the ECG dataset, we train ECG-Byte to compress the discretized ECG signals by iteratively merging the most frequent byte pairs into single tokens, following the BPE algorithm (Gage, 1994). The process starts by converting $\mathbf{X}_{\text{concat}}$ into a vector of token IDs derived from 8-bit byte values (stored as 32-bit unsigned integers) and initializing a vocabulary map (vocab) for string representations of bytes and a vocab_tokens map to encode bytes as singleton lists. IDs and vocab are initialized to cover the full byte range (0–255), mapping symbols in $\mathcal{A}$ to ASCII values (97–122), while reserving other byte values for unknown bytes. As merging proceeds, new tokens are assigned unique integer IDs starting from 256, acting as abstract labels for progressively larger token units. For each merge iteration, ECG-Byte calculates adjacent byte pair frequencies using a parallelized get_stats function, efficiently aggregating counts via a fold-and-reduce strategy. The most frequent pair is identified as the “best pair” to merge, and the merge function replaces occurrences of this pair in the ID vector with a new token ID, extending the vocabulary and updating vocab_tokens accordingly. This process repeats until the specified number of merges is reached or no pairs remain. The output includes the encoded ID vector, the extended vocabulary map, and a history of merge operations. Existing tokenizers, such as SentencePiece (Kudo and Richardson, 2018) or HuggingFace (Wolf et al., 2020), were not used due to their complexity and integration issues, which hindered interpretability. ECG-Byte, implemented in Rust for speed, provides a lightweight, flexible framework for representing ECG signals as discrete tokens while drawing inspiration from HuggingFace’s tokenizer (Wolf et al., 2020). We provide the detailed pseudocode of the training process in Algorithm 1 and pseudocode of the merge and get_stats functions in Appendix B.1.
Algorithm 1 Training Process for ECG-Byte
Input: Input $X_{\text{concat}}$, number of merges num_merges.
Output: Tuple containing final encoded IDs, vocabulary map, and merge history.
1:  ids $\leftarrow$ Convert $X_{\text{concat}}$ to a vector of $u32$
2:  vocab $\leftarrow$ Mappings from IDs 00 to $255$ as string
3:  vocab_tokens $\leftarrow$ Mappings from IDs 00 to $255$ to singleton lists
4:  merges $\leftarrow$ Empty list
5:  for $i\leftarrow 0$ to num_merges - 1 do
6:     pairs $\leftarrow$ get_stats(ids)
7:     if pairs is empty then
8:        break
9:     end if
10:     best_pair $\leftarrow$ pair in pairs with highest frequency
11:     if best_pair is not found then
12:        break
13:     end if
14:     new_id $\leftarrow 256+i$
15:     ids $\leftarrow$ merge(ids, best_pair, new_id)
16:     vocab[new_id] $\leftarrow$ concat(vocab[best_pair.0], vocab[best_pair.1])
17:     new_token $\leftarrow$ concat(vocab_tokens[best_pair.0], vocab_tokens[best_pair.1])
18:     vocab_tokens[new_id] $\leftarrow$ new_token
19:     Append (new_token, new_id) to merges
20:  end for
21:  return (ids, vocab, merges)